Project

General

Profile

Meeting 2014-04-16

Date: 16th April 2014
Time: 15:00 GMT
Attendance: E. Overton, M. A. Uchida

This week we just chatted about LED tests as we were the only ones here. We agreed that the PMT option seems the most sensible unless any major jolts are recorded on the accelerometer when we move the SS or if the outcome of the external waveguide QA shows that small cracks are a major problem. We will discuss it again in the next meeting with a wider audience.

Next meeting April 30th.

Agenda and Minutes

0. Actions

  • Create MICE note of channel mappings end-to-end through hardware, electronics and software - DA
  • Create comparison plots for channel mapping data - CH
  • Check with KL and AB if we want to proceed with new spare waveguides - MU
  • Place survey data on on tracker wiki - MU
  • Create GANTT chart for tracker integration in MICE hall - MU, CM
  • Prepare for LED testing, make connectors - GB, KLs

1. Software: AD, ES, CH, ST

  • Milestones - AD See Software Schedule.
  • Documentation - AD, ES, CH
    • Base documentation - AD
    • Digitisation - CH, ES
    • Pattern Recognition - AD
    • Track fit - ES
    • MC - CH
    • Geometry, CDB - CH
    • Noise - CH
  • Geometry, Calibrations, Configurations, CDB - CH
    • Get CMM geometry data in to CDB
    • Translate CDB geometry xml into MiceModules format Now able to pull down from CDB but still crashing when try to run sim with it
  • MC Noise - CH
  • ADC smearing needs fixing - CH
  • Final track fit ES off MICE for this month, pick up after that
    • Deal with -ve charge
  • Spacepoint efficiency study - C. Hunt, AD
  • Celeste effect - Resolved for muons, present for positrons, due to extremely low pt. Work ongoing.

2. DAQ/Calibration/Controls - DA, EO

2.1 MiniDAQ - DA

2.2 Channel Mappings

  • Check our channel mapping from end to end and make sure correct - DA, AD, ES, CH
  • Dig up the old documentation for the electronics channel mapping - DA

2.3 Controls

  • Create controls system - DA

2.4 Hall Probes - MU

  • Hall probe readout - ensure this is sorted before tracker placed in bore - PH, MU

3. LED: EO

  • Finish main flasher unit - EO, CMacW See EO presentation
  • Make documentation mapping LEDs to areas of tracker illuminated - EO
  • Ask KL, AB, DA about whether it would be helpful to fit LED into TOF to help our timing - EO

4. Tracker lab - AB, GB, CM

  • CM wants to use tracker lab to store tracker equipment currently spread around RAL
  • Move cryostat to go back to tracker lab. - CM

5. Installation

  • Build tent in R9 - GB Done, still issue of light tightness
  • Tracker integrity during transport from R9 to Hall
    • Very unlikely to be able to perform full LED readout test in R9
    • Add accelerometer when moving
  • Survey data
    • Need to obtain survey data from M. Leonova - MU Done
    • Store data sensibly (tracker wiki for now) - MU
    • Survey of phi needs to be done by John Pallin - GB
  • Finish installation manual - GB, MU
  • External waveguides
    • New connectors - GB Prices obtained, ~ £1600 for 10, £2500 for 25. Fibre will be more expensive, if we cannot use D0 fibre
    • Check with KL and AB if we want to proceed - MU
    • Need to QA the external waveguides somehow - MU
  • Create GANTT chart for tracker integration in MICE hall, and merge with main schedule - MU, CM

Hall probe alignment method etc agreed as surveying to tracker fiducial volume, line them up in same orientation, stick to number and positions agreed.
LED testing will commence in R9 after installation
Action Geoff to set ball in motion getting machinery necessary and starting Kevin making connectors
Craig to order new pulser box
all tbc by 7th April (optimistic deadline)
All has gone well ad to plan
top 1/2 tent up and lighting in and bore clean
Set for 5 day install push w/c/24th Marc

6. Services in the MICE hall: CM

  • Backing vacuum line system - CM

7. Mitigation of effect of magnetic field: CM, MU

Post box-esqu slit in PRY through which waveguides will feed has been extended downwards 1cm to crease. No increase in field but means top half of PRY can be removed without disturbing guides.

8. Step IV Readiness - MU

Work is ongoing, progress made but no results at this time.

9. AoB and summary of actions.

Phone Details

  • The four digit company id: 4673 (callers from outside the UK)
  • The pin number: 419357

Phone numbers:

UK (Imperial College --preferred) 0845 545 4460 English
Austria 0820 4000 1503 English
0820 4000 15 02 German
Belgium 070 35 99 45 English
070 35 99 94 Dutch
070 35 98 66 French
Canada 1 712 432 2832 English
1 712 432 2844 French
Germany 01803 001 178 English
01570 248 8303 English (mobile)
01803 001 177 German
01570 248 8101 German (mobile)
Ireland 0818 270 007 English
Italy 848 390 166 English
848 391 819 Italian
Netherlands 0870 001 901 English
0870 001 909 Dutch
Poland 0801 003 543 English
0801 003 533 Polish
South Africa 087 550 0375 English
Spain 902 885 318 English
902 881 200 Spanish
Sweden 0939 2066 400 English
0939 2066 300 Swedish
Switzerland 0848 560 190 English
0848 560 378 French
0848 560 347 German
0848 560 385 Italian
United Kingdom (PowWowNow) 0844 4 73 73 73 English
United States 1 415 363 0833 English